OpenStudy (anonymous):

Find the dimensions of the rectangle with largest area that can be inscribed in a semicircle of radius 2 inches.

OpenStudy (anonymous):

radius = 2 A= lw = 2xy cos beta =x/r sine beta = y/r 2cos beta =x 2sin beta =y so ...A= 2x(4cosbetasinebeta) Area equals to 8 time cosine beta time sine beta so just plug in 15 , 30, 45, 60, 75 degree to the formula...then u noticed that at degree 45 the area is four which is the largest area after that. u know perimeter is 2x+2y since area eaquals 2xy then y equals to 4/2x then plug it in the perimeter equatin then you would get 2x +4/x d/dy 2x+4/x then you would get radical 2 for y